OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4) answer has to be a certain way..

OpenStudy (anonymous):

-->fill in spaces.. ___x___y_________ @some1

OpenStudy (anonymous):

what does it mean "fill in spaces"?

OpenStudy (anonymous):

ohh the answer would be 3x+y=

OpenStudy (anonymous):

@some1 can u help?

OpenStudy (anonymous):

what do i do after 24+(-4)=b

OpenStudy (anonymous):

is it b=28?

OpenStudy (anonymous):

did you find the slope?

OpenStudy (anonymous):

oh yes, you have it

OpenStudy (anonymous):

okai so it seems to me you did it right. exceppppt 24-4 is not 28 :P

OpenStudy (anonymous):

so 20 right?

OpenStudy (anonymous):

so its 3x+y=20

OpenStudy (anonymous):

yes \[3x+y=20\] or \[3x+y-20=0\] which ever form they want

OpenStudy (anonymous):

yep. that's right shell =)

OpenStudy (anonymous):

woohew, k thanks:D

OpenStudy (anonymous):

np :) u're getting better!

OpenStudy (anonymous):

because of your smartness -.- :D

OpenStudy (anonymous):

nouuuu it's cuz YOU'RE smart.

OpenStudy (anonymous):

lol sure...

OpenStudy (anonymous):

hey.. is this --> M(0, 6), N(6, 0).... slope... = 1.. or -1?

OpenStudy (anonymous):

just 1 .. right?

OpenStudy (anonymous):

0-6/6-0=-1.

OpenStudy (anonymous):

k:D

OpenStudy (anonymous):

so it would be x+y=-6?

OpenStudy (anonymous):

yush! =)

OpenStudy (anonymous):

:D yaaaaaaaaaaaaaaaaaaaay:D